Find the standard deviation of the following distribution

X -9 -6 -3 0 3
P(x) 0.2 0.1 0.2 0.3 0.2
Answer #1

Solution,

X P(x) X * P(x) X2 * P(x)
-9 0.2 -1.8 16.2
-6 0.1 -0.6 3.6
-3 0.2 -0.6 1.8
0 0.3 0 0
3 0.2 0.6 1.8
\sum 1.00 -2.4 23.4

a) mean = u = \sum x * p(x) = -2.4

b) standard deviation = \sigma = \sqrt{} \sum x2 * p(x) - (u)2

standard deviation = \sigma = \sqrt{} 23.4 - ( -2.4)2

standard deviation = \sigma = \sqrt{} 23.4 - 5.76

standard deviation = \sigma = \sqrt{} 17.64

standard deviation = \sigma = 4.2